A financial analyst has found out that, policyholders are four times as likely to fi le two
claims as to fi le four claims. If the number of claims led has a Poisson distribution,
what is the expected value of the number of claims led?
1
Expert's answer
2020-06-04T19:37:24-0400

Let "X=" the number of claims: "X\sim Po(\lambda)"


"P(X=x)={e^{-\lambda}\cdot \lambda^x\over x!}, \lambda>0"

Given "P(X=2)=4P(X=4)"


"P(X=2)={e^{-\lambda}\cdot \lambda^2\over 2!}=4P(X=4)=4{e^{-\lambda}\cdot \lambda^4\over 4!}"

We want to find "\lambda" for the Poisson RV


"{ \lambda^2\over 2}=4\cdot{ \lambda^4\over 24}, \lambda>0"

"\lambda^2=3, \lambda>0"

So "\lambda=\sqrt{3}."

The expected value of the number of claims led is "\sqrt{3}."